# Single-shot and measurement-based quantum error correction via fault complexes

### v1

@@ Abstract
Photonics provides a viable path to a scalable fault-tolerant quantum computer. The natural framework for this platform is measurement-based quantum computation, where fault-tolerant graph states supersede traditional quantum error-correcting codes. However, the existing formalism for foliation—the construction of fault-tolerant graph states—does not reveal how certain properties, such as single-shot error correction, manifest in the measurement-based setting. We introduce the fault complex, a representation of dynamic quantum error correction protocols particularly well-suited to describe foliation. Our approach enables precise computation of fault tolerance properties of foliated codes and provides insights into circuit-based quantum computation. Analyzing the fault complex leads to improved thresholds for three- and four-dimensional toric codes, a generalization of stability experiments, and the existence of single-shot lattice surgery with higher-dimensional topological codes.
@@
Introduction.— Certain scalable platforms for fault-tolerant quantum computers, namely photonics [1, 2, 3, 4], are particularly well-suited to the model of measurement-based quantum computation (MBQC) [5]. In this paradigm, in contrast to circuit-based quantum computation (CBQC), the central object is not a quantum error-correcting (QEC) code but a fault-tolerant graph state (FTGS). Various methods for constructing an FTGS exist [6, 7, 8], with the concept of foliation [9, 10] offering a prescription for any Calderbank-Shor-Steane (CSS) code [11, 12]. Here, we introduce the fault complex, a representation of faults in a dynamic QEC protocol rather than a static QEC code. We focus on fault complexes obtained via foliation, which we recast in the language of homology as a tensor product [13, 14] between a CSS code and a repetition code. This mathematical machinery allows us to exactly compute the properties of the QEC protocol, such as the fault distance, starting from any input CSS code. Our formalism is based on the representation of CSS codes as chain complexes [15, 16] and is related to distance balancing [17, 18]. Fault complexes also have an interpretation as repeated rounds of error correction over time in CBQC, allowing us to apply insights from our formalism to that domain. Analysis of the fault complex enables us to understand the decoding of single-shot codes in MBQC and achieve improved error thresholds for three-dimensional (3D) and 4D toric codes [19, 20, 21, 22]. Through the explicit calculation of the homology groups of the fault complex, we generalize the notion of stability experiments [23], which in turn allows us to conclude that single-shot lattice surgery is possible in higher-dimensional topological codes.
@@
Background and Notation.— An $[n,k,d]$ binary linear code $C$ forms a $k$-dimensional subspace within the $n$-dimensional vector space over $\mathbb{F}_{2}$. Such a code can be defined by a parity-check matrix (PCM) $H$, where the codewords are the elements of $\ker H$. The distance $d$ is the minimum Hamming weight of any nonzero codeword. We use $e_{i}$ to denote the $i^{\text{th}}$ unit vector and $[M]_{j}$ to denote the $j^{\text{th}}$ row of a matrix $M$. We write $\mathbf{0}$ and $\mathbf{1}$ for the all-zero and all-one vectors, respectively.
@@
A stabilizer code [24] is the quantum analogue of a linear code and is defined by an abelian subgroup, $\mathcal{S}$, of the Pauli group with $-I\notin\mathcal{S}$. The codespace is the $+1$ eigenspace of $\mathcal{S}$; logical operators commute with $\mathcal{S}$ but are not themselves in $\mathcal{S}$. The notation $\llbracket n,k,d\rrbracket$ describes a code in which the stabilizer group is generated by $m=n-k$ independent generators, which are themselves elements of the $n$-qubit Pauli group $\mathcal{P}_{n}$. The number of encoded qubits is $k$ and the distance $d$ is determined by the minimum weight non-trivial logical operator. A stabilizer code admitting a set of stabilizer generators that are each either $X$-type or $Z$-type operators is referred to as a CSS code [11, 12]. These codes can be described by two classical binary linear codes with PCMs $H_{X}$ and $H_{Z}$ representing stabilizer generators as tensor products of $X$ and $I$, and $Z$ and $I$, respectively. Commutativity of $X$- and $Z$-type generators is expressed through the condition $H_{Z}^{T}H_{X}=0$. Error correction proceeds via the measurement of the stabilizer generators, yielding the syndrome $s$, which is the list of stabilizer eigenvalues. This allows the detection and correction of Pauli errors that anticommute with stabilizer generators; a classical decoder is used to infer recovery operators given a syndrome.
@@
with the composition fulfilling $\partial_{i}^{C}\partial_{i+1}^{C}=0$. We suppress superscripts if no distinction is necessary. Elements of the kernel $Z_{i}(\mathcal{C})\coloneqq\ker\partial_{i}$ are referred to as $i$-cycles, and elements of the image $B_{i}(\mathcal{C})\coloneqq\operatorname{im}\partial_{i+1}$ as $i$-boundaries. The quotient $H_{i}(\mathcal{C})\coloneqq Z_{i}(\mathcal{C})/B_{i}(\mathcal{C})=\ker\partial_{i}/\operatorname{im}\partial_{i+1}$ is called the $i^{\text{th}}$ homology group of $\mathcal{C}$. Associated with $\mathcal{C}$ is also a cochain complex, with coboundary operators $\delta^{i}:C_{i}\rightarrow C_{i+1}$ defined as $\delta^{i}=\partial_{i+1}^{T}$:
$\mathcal{C}^{T}=\{0\}\stackrel{{\scriptstyle\delta^{-1}}}{{\rightarrow}}C_{0}\stackrel{{\scriptstyle\delta^{0}}}{{\rightarrow}}C_{1}\dots\stackrel{{\scriptstyle\delta^{n-1}}}{{\rightarrow}}C_{n}\stackrel{{\scriptstyle\delta^{n}}}{{\rightarrow}}\{0\}.$ (2)
Elements of $Z^{i}\coloneqq\ker\delta^{i}$ and $B^{i}\coloneqq\operatorname{im}\delta^{i-1}$ are referred to as cocycles and coboundaries, respectively, and $H^{i}\coloneqq\ker\delta^{i}/\operatorname{im}\delta^{i-1}$ is the $i^{\text{th}}$ cohomology group.
Each length-1 chain complex describes an $[n,k,d]$ linear code with boundary map $\partial_{1}=H$ from the vector space $\mathbb{F}_{2}^{n}$ to the space of syndromes $\mathbb{F}_{2}^{r}$ with $r\geq n-k$. Similarly, a CSS code, $\mathcal{C}$, can represented by a length-2 chain (sub)complex
@@
Fault Complexes.— We define a fault complex to be a length-3 chain (sub)complex $\mathcal{F}$
@@
Recently, there have been many proposals for formalizing fault-tolerant protocols [7, 8, 25, 26, 27, 28, 29, 30, 31, 32], some of which bear some resemblance to fault complexes. In particular, our definition of a fault complex builds on the definition of a fault-tolerant cluster state in [8]. We also note that fault complexes are distinct from the fault-tolerant complexes of [29], which are defined geometrically and are limited to topological codes. Fault complexes have an interpretation in CBQC where the faults can represent both qubit, gate, and measurement errors. In this interpretation, $\partial_{i+1}$ is related to the gauge group of the spacetime code; see Appendix F for a further discussion. Here, we concentrate on the interpretation of fault complexes in MBQC, where they give a useful representation of foliated codes.
Fault Complexes of Foliated Codes.— A common way to obtain an FTGS from a CSS code is through foliation [9]. In this procedure, the $X$ and $Z$ Tanner graphs of the code are interpreted as alternating layers of the FTGS, and data qubits between adjacent layers are entangled with each other; see Fig. 1(a-d). The computation is then performed by a sequence of adaptive local measurements on the FTGS. We now show that foliation can be algebraically formulated as the tensor (or hypergraph [13, 14]) product of a base CSS code with a repetition code, yielding a fault complex. As a result, the detectors, logical correlations/errors, and fault distances of the fault complex can be directly calculated from the base code.
Figure 1: Foliation of the surface code viewed as a fault complex. (a) The distance-3 surface code Tanner graph with circles representing qubits, squares with appropriate labels representing $X$ and $Z$ checks, and dashed and solid lines representing the connectivity of these checks, respectively. (b) and (c) Hypergraph product of (a) with a repetition code check and bit node, respectively. This yields two types of fault locations, teal and yellow, along with primal (red) and dual (purple) detectors. Dashed and solid lines indicate boundary maps of the fault complex. For simplicity, we omit out-of-plane connections. (d) Unit cell of the fault complex obtained by stacking alternating layers of (b) and (c). A dual detector (purple) is formed by the parity of six dual fault locations (yellow). Vertical red dashed lines are omitted for clarity. (e) Mathematical structure of the fault complex and its relation to (b)-(d): colored circles and squares are associated with chain complex vector spaces and lines are associated with boundary maps.
Let $\mathcal{C}$ be the chain complex describing a CSS code and let $\mathcal{R}$ be a length-1 chain complex describing a binary linear code. We obtain the fault complex as the product complex =×$\mathcal{F}=\mathcal{R}\times\mathcal{C}$. The fault complex has spaces
$F_{j}=\bigoplus_{\ell+m=j}R_{\ell}\otimes C_{m}$ (5)
where $R_{\ell}$ and $C_{m}$ are the $\ell^{\text{th}}$ and $m^{\text{th}}$ spaces of $\mathcal{R}$ and $\mathcal{C}$, respectively. The boundary operators of the fault complex are
where $\partial^{C}_{j}$ is the $j^{\text{th}}$ boundary operator of $\mathcal{C}$, and $R$ is the $r\times c$ PCM of the binary linear code. See Fig. 1(e) for a diagrammatic representation of the fault complex vector space and boundary maps. From the Künneth formula, we obtain the number of primal and correlations
@@
One can show [14, and Appendix B] that the primal and dual fault distances of $\mathcal{F}$ are given by
@@
Example: CSS codes.— Suppose that $\mathcal{R}$ describes a repetition code with PCM $R$ and distance $\delta$, and that $\mathcal{C}$ is a length-2 chain complex describing a CSS code with PCMs $H_{X}$ and $H_{Z}$. Further, we assume that $R$, $H_{X}$, and $H_{Z}$ are full rank. The nontrivial boundary operators, $\partial_{j},j\in\{1,2,3\}$, of the fault complex =×$\mathcal{F}=\mathcal{R}\times\mathcal{C}$ can be explicitly obtained from Eq. 8. In particular, the rows of $\partial_{1}$ describe the support of primal detectors; they are all of the form $(e_{\alpha}\otimes[H_{X}]_{\beta}|[R]_{\alpha}\otimes e_{\beta})$, with an analogous form for the columns of $\partial_{3}$. These are exactly the foliated stabilizers of [9]; see Fig. 1(d), which illustrates a dual detector. Each detector (row) above is triggered by a fault on one of the qubits participating in the check $[H_{X}]_{\beta}$, or by a syndrome fault in the previous or next layer. The logical correlations are recovered from the homology group, that is, for example,
@@
where $\cong$ denotes isomorphic to and $\ell_{Z}$ is a logical $Z$ operator of the base code. See Appendix B for a detailed derivation. This operator is a dual logical correlation; it acts as a copy of the logical $Z$ operator on the qubits in the $R_{1}\otimes C_{1}$ block of the $F_{1}$ space. The dual logical errors are given by
@@
where $\ell_{X}$ is a logical $X$ operator of the base code. We observe that this operator acts as a logical $X$ operator on one of the factors in the $R_{1}\otimes C_{1}$ portion of the $F_{1}$ space. Applying Eq. 9, we find that the (dual) fault distance of $\mathcal{F}$ is $d_{2}=d_{X}$, where $d_{X}$ is the $X$-distance of the base code.
For our choice of $\mathcal{R}$, the homology groups $H_{0}(\mathcal{R})$ and $H^{0}(\mathcal{R})$ are trivial, so the fault complex lacks primal logical correlations arising from the base code’s logical $Z$. This may seem surprising, given that the foliated surface code can be used to prepare an encoded Bell state on the two boundaries [33]. There is no contradiction, however: the interpretation of fault complexes as foliated codes implicitly assumes that all qubits are measured in the $X$ basis during the protocol, and therefore that the encoded Bell state of [33] is measured destructively. The analogous CBQC interpretation is of a memory experiment with logical state preparation and readout both performed in the $Z$ basis. In addition, we note that all the logical correlations can be recovered by considering an alternative repetition code PCM (see Appendix B).
Stability experiments.— Now suppose that $H_{X}$ of the base code $\mathcal{C}$ is rank deficient. The fault complex =×$\mathcal{F}=\mathcal{R}\times\mathcal{C}$ now has additional logical correlations and errors coming, respectively, from the following homology groups
@@
Here, $g$ represents any subset of $H_{X}$ rows that sum to zero and $h$ is any vector in $C_{0}$ that is not the syndrome of a $Z$-type error. An example of such a code is a 2D surface code with four smooth boundaries; in this case we have $g=\bf 1$ (the product of all $X$ stabilizers is $I$) and $h=(1,0,\ldots,0)$ (all valid syndromes have even weight). Note that $g$ is exactly the global spacelike invariant considered in stability experiments [23], which can be used to estimate the expected logical error when performing lattice surgery [34, 35], a method for performing logical operations via measurements. Using Eq. 9 we find that the primal distance of $\mathcal{F}$ is $d_{1}=\delta$, i.e., equal the distance of the repetition code (or the number of CBQC measurement rounds cf. [23]). A 2D surface code with four smooth boundaries encodes no logical qubits, but if instead we let $\mathcal{C}$ be the 2D toric code (with periodic boundaries) then our formalism shows that the fault complex =×$\mathcal{F}=\mathcal{R}\times\mathcal{C}$ can be used to perform a combined memory and stability experiment.
@@
As a prototypical example of a code with metachecks, let $\mathcal{C}$ represent the 3D toric code [19, 20] defined on an $L\times L\times L$ cubic tiling, which has single-shot QEC for $Z$ errors. We focus on this code for ease of presentation, but our results also hold for true single-shot CSS codes such as the 4D toric code. We assign the metacheck matrix $M_{X}=\partial_{1}^{C}$ and PCMs $H_{X}=\partial_{2}^{C}$ and $H_{Z}=(\partial_{3}^{C})^{T}$. The primal detectors of the fault complex =×$\mathcal{F}=\mathcal{R}\times\mathcal{C}$ are given by $\partial_{2}$, which now contains the metachecks; see Appendix C for the explicit expression. We note that $\mathcal{F}$ is effectively a length-3 chain complex exactly because ‘data’ and ‘measurement’ faults are treated on an equal footing.
@@
The expression for the primal logical correlations and errors of $\mathcal{F}$ is the same as Eq. 12, but now with $g\in\ker H_{X}^{T}/\operatorname{im}M_{X}^{T}$ and $h\in\ker M_{X}/\operatorname{im}H_{X}$. As in the previous example, $g$ is the logical correlation relevant for stability experiments and lattice surgery, and $h$ is the logical error that can disrupt this correlation. In the 3D toric code, there are three independent choices of $g$ given by all edges cutting through one of the independent 2D planes of the tiling. The corresponding $h$ vectors correspond to non-contractible chains of edges along one of the coordinate axes of the tiling. As a result, all choices of $h$ have extensive weight, i.e., $|h|\geq L$. Thus, even for a fault complex formed using a constant-length repetition code, a macroscopic number of faults (${\sim}L$) is necessary to disrupt the $g$ logical correlations. This is captured by the primal fault distance of $\mathcal{F}$, $d_{1}=\delta L$, where $\delta=d_{1}(\mathcal{R})$ is the distance of the repetition code. Therefore, higher dimensional topological codes such as the 3D and 4D toric codes are compatible with single-shot lattice surgery, though at the cost of reduced performance. Counter-intuitively, the full distance can be restored in those cases by choosing $\mathcal{R}$ such that $\delta=L=O(\sqrt{d})$, (where $d$ is the distance of the code), or equivalently by performing $O(\sqrt{d})$ rounds of stabilizer measurement in the lattice surgery protocol. This should be contrasted with the 2D toric code case where the number of measurement rounds must be $O(d)$. Thus, we expect that a fault-tolerant quantum computing architecture based on the 4D toric code would have an asymptotic spacetime overhead reduction when compared to the standard 2D toric code architecture.
@@
We address the case of single-shot codes with rank-deficient PCMs without metachecks in Appendix B. We note that there exist single-shot quantum codes with full-rank PCMs [38, 39], to which our arguments above cannot be applied, and the existence of single-shot lattice surgery protocols for these codes remains an open question. We conjecture that codes enabling single-shot lattice surgery without performance degradation can be constructed from the balanced product of two good qLDPC codes [40, 41, 42, 43, 44].
@@
Improved decoding of single-shot codes.— Estimating the error threshold of a single-shot code requires simulating multiple rounds of noisy syndrome measurement until the threshold has converged [45]. Since in MBQC there always exist detectors spanning multiple rounds, decoding must proceed using an overlapping window decoder [19, 46, 47, 48, 49], which is naturally defined for a fault complex.
Figure 2: Sustainable thresholds for 3D (top row) and 4D (bottom row) toric codes under two noise models: phenomenological Pauli noise (left column) and a noise model for a photonic GKP-based architecture (right column). Markers represent the threshold for a $(w,1)$-overlapping window decoder as a function of noisy syndrome rounds. The black dashed line shows the threshold from the optimal window choice $w=L$; see Appendix D for details. In the 3D case we only consider primal faults, whereas in the 4D case we consider both primal and dual faults.

### v2

@@ Abstract
Photonics provides a viable path to a scalable fault-tolerant quantum computer. The natural framework for this platform is measurement-based quantum computation, where fault-tolerant graph states supersede traditional quantum error-correcting codes. However, the existing formalism for foliation—the construction of fault-tolerant graph states—does not reveal how certain properties, such as single-shot error correction, manifest in the measurement-based setting. We introduce the fault complex, a representation of dynamic quantum error correction protocols particularly well-suited to describe foliation. Our approach enables precise computation of fault tolerance properties of foliated codes and provides insights into circuit-based quantum computation. Analyzing the fault complex yields improved thresholds for three- and four-dimensional toric codes, a generalization of stability experiments, and the existence of single-shot lattice surgery with higher-dimensional topological codes.
@@
Introduction.— Photonic platforms for quantum computing [1, 2, 3, 4, 5, 6] are well-suited to measurement-based quantum computing (MBQC) [7]. In this paradigm, in contrast to circuit-based quantum computation (CBQC), the central object is not a quantum error-correcting (QEC) code but a fault-tolerant graph state (FTGS). Various methods for constructing FTGS exist [8, 9, 10], with the concept of foliation [11, 12] offering a prescription for any Calderbank-Shor-Steane (CSS) code [13, 14]. Here, we introduce the fault complex, a representation of faults in a dynamic QEC protocol (rather than a static QEC code) formulated in the language of homology and chain complexes [15, 16, 17, 18]. We focus on fault complexes obtained via foliation, which we recast in the language of homology as a tensor product [19, 20] between a CSS code and a repetition code. This formalism allows us to easily calculate properties like the fault distance, and also applies to CBQC where it represents repeated rounds of error correction. Analysis of the fault complex enables us to understand the decoding of single-shot codes in MBQC and achieve improved error thresholds for three-dimensional (3D) and 4D toric codes [21, 22, 23, 24]. Through the explicit calculation of the homology groups of the fault complex, we generalize the notion of stability experiments [25], which in turn allows us to conclude that single-shot lattice surgery is possible in higher-dimensional topological codes. The fault complex provides a formal language for fault-tolerant protocols, similar to how homological descriptions serve CSS codes, laying the groundwork for future advances in fault-tolerant quantum computation.
@@
Background and Notation.— An $[n,k,d]$ binary linear code $C$ forms a $k$-dimensional subspace within the $n$-dimensional vector space over $\mathbb{F}_{2}$. Such a code can be defined by a parity-check matrix (PCM) $H$, where the codewords are the elements of $\ker H$. The distance $d$ is the minimum Hamming weight of any nonzero codeword. We use $e_{i}$ to denote the $i^{\text{th}}$ unit vector and $M]_{j}$ to denote the $j^{\text{th}}$ row of a matrix $M$. We write $\mathbf{0}$ and $\mathbf{1}$ for the all-zero and all-one vectors, respectively.
@@
A stabilizer code [26] is the quantum analogue of a linear code and is defined by an abelian subgroup, $\mathcal{S}$, of the Pauli group with $I\notin\mathcal{S}$. The codespace is the +1+1 eigenspace of $\mathcal{S}$; logical operators commute with $\mathcal{S}$ but are not themselves in $\mathcal{S}$. The notation $\llbracket n,k,d\rrbracket$ describes a code in which the stabilizer group is generated by m=nkm=n-k independent generators, which are themselves elements of the $n$-qubit Pauli group $\mathcal{P}_{n}$. The number of encoded qubits is $k$ and the distance $d$ is determined by the minimum weight non-trivial logical operator. A stabilizer code admitting a set of stabilizer generators that are each either $X$-type or $Z$-type operators is referred to as a CSS code [13, 14]. These codes can be described by two classical binary linear codes with PCMs $H_{X}$ and $H_{Z}$ representing stabilizer generators as tensor products of $X$ and $I$, and $Z$ and $I$, respectively. Commutativity of $X$- and $Z$-type generators is expressed through the condition $H_{Z}^{T}H_{X}=0$. Error correction proceeds via the measurement of the stabilizer generators, yielding the syndrome $s$, which is the list of stabilizer eigenvalues. This allows the detection and correction of Pauli errors that anticommute with stabilizer generators; a classical decoder is used to infer recovery operators given a syndrome.
@@
with the composition fulfilling $\partial_{i}^{C}\partial_{i+1}^{C}=0$. We suppress superscripts if no distinction is necessary. The quotient $H_{i}(\mathcal{C})\coloneqq\ker\partial_{i}/\operatorname{im}\partial_{i+1}$ is called the $i^{\text{th}}$ homology group of $\mathcal{C}$. Associated with $\mathcal{C}$ is also a cochain complex, with coboundary operators $\delta^{i}:C_{i}\rightarrow C_{i+1}$ defined as $\delta^{i}=\partial_{i+1}^{T}$:
$\mathcal{C}^{T}=\{0\}\stackrel{{\scriptstyle\delta^{-1}}}{{\rightarrow}}C_{0}\stackrel{{\scriptstyle\delta^{0}}}{{\rightarrow}}C_{1}\dots\stackrel{{\scriptstyle\delta^{n-1}}}{{\rightarrow}}C_{n}\stackrel{{\scriptstyle\delta^{n}}}{{\rightarrow}}\{0\},$ (2)
with cohomology groups $H^{i}\coloneqq\ker\delta^{i}/\operatorname{im}\delta^{i-1}$.
Each length-1 chain complex describes an $[n,k,d]$ linear code with boundary map $\partial_{1}=H$ from $\mathbb{F}_{2}^{n}$ to the space of syndromes $\mathbb{F}_{2}^{r}$ with $r\geq n-k$. Similarly, a CSS code, $\mathcal{C}$, can represented by a length-2 chain (sub)complex
@@
Figure 1: Foliation of the surface code viewed as a fault complex. (a) The distance-3 surface code Tanner graph with circles representing qubits, squares representing $X$ and $Z$ checks, and dashed and solid lines representing the connectivity of these checks, respectively. (b) and (c) Hypergraph product of (a) with a repetition code check and bit node, respectively. This yields two types of fault locations, teal and yellow, along with primal (red) and dual (purple) detectors. Dashed and solid lines indicate boundary maps of the fault complex. For simplicity, we omit out-of-plane connections. (d) Unit cell of the fault complex obtained by stacking alternating layers of (b) and (c). A dual detector (purple) is formed by the parity of six dual fault locations (yellow). Vertical red dashed lines are omitted for clarity. (e) Mathematical structure of the fault complex and its relation to (b)-(d): colored circles and squares represent chain complex vector spaces and lines represent boundary maps.
Motivating example.— A common way to obtain an FTGS from a CSS code is via foliation [11]. In this procedure, the $X$ and $Z$ Tanner graphs of the code are interpreted as alternating layers of the FTGS, and data qubits between adjacent layers are entangled with each other; see Fig. 1(a-d). The computation is then performed by a sequence of adaptive local measurements on the FTGS. We observe that foliation can be algebraically formulated as the tensor (or hypergraph [19, 20]) product of a base CSS code with a repetition code  [^1] We note that a similar approach has recently explored in the context of condensed matter physics [27, 28]; see Fig. 1e. Denote the resulting complex as $\mathcal{F}=\mathcal{C}\times\mathcal{R}$, where $\mathcal{C}$ is a length-2 chain complex describing a CSS code and $\mathcal{R}$ is a chain complex describing a repetition code. The spaces of $\mathcal{F}$ are
$F_{j}=\bigoplus_{\ell+m=j}R_{\ell}\otimes C_{m}$ (4)
where $R_{\ell}$ and $C_{m}$ are the $\ell^{\text{th}}$ and $m^{\text{th}}$ spaces of $\mathcal{R}$ and $\mathcal{C}$, respectively. The boundary operators of $\mathcal{F}$ are
where $\partial^{C}_{j}$ is the $j^{\text{th}}$ boundary operator of $\mathcal{C}$, and $R$ is the $r\times c$ PCM of the binary linear code. One can observe that $\partial_{2}$ is the bi-adjacency matrix of the (bipartite) FTGS, whereas $\partial_{1}$ and $\partial_{3}^{T}$ describe the detectors (or foliated stabilizers) of the FTGS. For example, the rows of $\partial_{1}$ are of the form $(e_{\alpha}\otimes[H_{X}]_{\beta}|[R]_{\alpha}\otimes e_{\beta})$, with an analogous form for the columns of $\partial_{3}$. These are exactly the foliated stabilizers of [11]; see Fig. 1(d), which illustrates a dual detector. Each detector (row) above is triggered by a fault on one of the qubits participating in the check $H_{X}]_{\beta}$, or by a syndrome fault in an adjacent layer.
Fault complexes.— Motivated by the previous example, we define a fault complex to be a length-3 chain (sub)complex $\mathcal{F}$
@@
Recently, there have been many proposals for formalizing fault-tolerant protocols [9, 10, 29, 30, 31, 32, 33, 34, 35, 36], some of which resemble our fault complexes. In particular, our definition of a fault complex builds on the definition of a fault-tolerant cluster state in [10]. We also note that fault complexes are distinct from the fault-tolerant complexes of [33], which are defined geometrically and are limited to topological codes. Fault complexes have an interpretation in CBQC where the faults can represent both qubit, gate, and measurement errors. In this interpretation, $\partial_{i+1}$ is related to the gauge group of the spacetime code; see [37] for further discussion. Here, we concentrate on the interpretation of fault complexes in MBQC.
Foliated CSS codes.— We now return to our motivating example of the fault complex $\mathcal{F}=\mathcal{C}\times\mathcal{R}$, where $\mathcal{C}$ represents a CSS code and $\mathcal{R}$ represents a repetition code. Suppose that $R$, $H_{X}$, and $H_{Z}$ are full rank. From the Künneth formula, we obtain the number of primal and dual correlations
@@
One can show [20, 37] that the primal and dual fault distances of $\mathcal{F}$ are given by
@@
The logical correlations of $\mathcal{F}$ are recovered from the homology group, that is, for example,
@@
where $\ell_{Z}$ is a logical $Z$ operator of the base code. See [37] for a detailed derivation. This operator is a dual logical correlation; it acts as a copy of the logical $Z$ operator on the qubits in the $R_{1}\otimes C_{1}$ block of the $F_{1}$ space. The dual logical errors are given by
@@
where $\ell_{X}$ is a logical $X$ operator of the base code. We observe that this operator acts as a logical $X$ operator on one of the factors in the $R_{1}\otimes C_{1}$ portion of the $F_{1}$ space. Applying Eq. 9, we find that the (dual) fault distance of $\mathcal{F}$ is $d_{2}=d_{X}$, the $X$-distance of the base code.
For our choice of $\mathcal{R}$, the homology groups $H_{0}(\mathcal{R})$ and $H^{0}(\mathcal{R})$ are trivial, so the fault complex lacks primal logical correlations arising from the base code’s logical $X$. This may seem surprising, given that the foliated surface code can be used to prepare an encoded Bell state on the two boundaries [38]. There is no contradiction, however: the interpretation of fault complexes as foliated codes implicitly assumes that all qubits are measured in the $X$ basis during the protocol, and therefore that the encoded Bell state of [38] is measured destructively. The analogous CBQC interpretation is of a memory experiment with logical state preparation and readout both performed in the $Z$ basis. In addition, we note that all the logical correlations can be recovered by considering an alternative repetition code PCM; see [37].
Stability experiments.— Lattice surgery [39] is the leading technique for performing logical operations on topological codes [40, 41], and stability experiments [25] estimate the logical error during a lattice surgery operation. Such experiments test our ability to accurately measure the product of many stabilizer generators.
Returning to our foliated CSS code example, suppose that $H_{X}$ of the base code $\mathcal{C}$ is rank deficient. The fault complex $\mathcal{F}=\mathcal{R}\times\mathcal{C}$ now has additional logical correlations and errors coming, respectively, from the following homology groups
@@
Here, $g$ represents any subset of $H_{X}$ rows that sum to zero and $h$ is any vector in $C_{0}$ that is not the syndrome of a $Z$-type error. An example of such a code is a 2D surface code with four smooth boundaries; in this case we have $g=\bf 1$ (the product of all $X$ stabilizers is $I$) and $h=(1,0,\ldots,0)$ (all valid syndromes have even weight). The logical correlation $g$ is exactly the product of stabilizer generators considered in stability experiments. Using Eq. 9 we find that the primal distance of $\mathcal{F}$ is $d_{1}=\delta$, i.e., equal the distance of the repetition code (or the number of CBQC measurement rounds cf. [25]). A 2D surface code with four smooth boundaries encodes no logical qubits, but if instead we let $\mathcal{C}$ be the 2D toric code (with periodic boundaries) then our formalism shows that the fault complex $\mathcal{F}=\mathcal{R}\times\mathcal{C}$ can be used to perform a combined memory and stability experiment; see [37].
@@
As a prototypical example of a code with metachecks, let $\mathcal{C}$ represent the 3D toric code [21, 22] defined on an $L\times L\times L$ cubic tiling, which has single-shot QEC for $Z$ errors. We focus on this code for ease of presentation, but our results also hold for true single-shot CSS codes such as the 4D toric code. We assign the metacheck matrix $M_{X}=\partial_{1}^{C}$ and PCMs $H_{X}=\partial_{2}^{C}$ and $H_{Z}=(\partial_{3}^{C})^{T}$. The primal detectors of the fault complex $\mathcal{F}=\mathcal{R}\times\mathcal{C}$ are given by $\partial_{2}$, which now contains the metachecks; see [37] for the explicit expression. We note that $\mathcal{F}$ is effectively a length-3 chain complex exactly because ‘data’ and ‘measurement’ faults are treated on an equal footing, see [37] for an extended discussion.
@@
The expression for the primal logical correlations and errors of $\mathcal{F}$ is the same as Eq. 12, but now with $g\in\ker H_{X}^{T}/\operatorname{im}M_{X}^{T}$ and $h\in\ker M_{X}/\operatorname{im}H_{X}$. As in the previous example, $g$ is the logical correlation relevant for stability experiments and lattice surgery, and $h$ is the logical error that can disrupt this correlation. In the 3D toric code, there are three independent choices of $g$ given by all edges cutting through one of the independent 2D planes of the tiling. The corresponding $h$ vectors correspond to non-contractible chains of edges along one of the coordinate axes of the tiling. As a result, all choices of $h$ have extensive weight, i.e., $|h|\geq L$. Thus, even for a fault complex formed using a constant-length repetition code, a macroscopic number of faults (${\sim}L$) is necessary to disrupt the $g$ logical correlations. This is captured by the primal fault distance of $\mathcal{F}$, $d_{1}=\delta L$, where $\delta=d_{1}(\mathcal{R})$ is the distance of the repetition code. Therefore, higher-dimensional topological codes such as the 3D and 4D toric codes are compatible with single-shot lattice surgery, though at the cost of reduced performance. The full distance can be restored by choosing $\mathcal{R}$ such that $\delta=L=O(\sqrt{d})$, (where $d$ is the distance of the code), or equivalently by performing $O(\sqrt{d})$ rounds of stabilizer measurement in the lattice surgery protocol. This should be contrasted with the 2D toric code case, where the number of measurement rounds must be $O(d)$. Thus, we expect that a fault-tolerant quantum computing architecture based on the 4D toric code would have an asymptotic spacetime overhead reduction when compared to the standard 2D toric code architecture.
@@
We address the case of single-shot codes with rank-deficient PCMs without metachecks in [37]. We note that there exist single-shot quantum codes with full-rank PCMs [44, 45], to which our arguments above cannot be applied, and the existence of single-shot lattice surgery protocols for these codes remains an open question. We conjecture that codes enabling single-shot lattice surgery without performance degradation can be constructed from the balanced product of two good qLDPC codes [46, 47, 48, 49, 50].
@@
Improved decoding of single-shot codes.— Estimating the error threshold of a single-shot code requires simulating multiple rounds of noisy syndrome measurement until the threshold has converged [51]. Since in MBQC there always exist detectors spanning multiple rounds, decoding must proceed using an overlapping window decoder [21, 52, 53, 54, 55], which is naturally defined for a fault complex. The fault complex framework provides a systematic guide for constructing efficient simulations and evaluating the performance of these protocols.